OpenStudy (anonymous):

A cone has a volume of 12 cubic inches. What is the volume of a cylinder that the cone fits exactly inside of? (1 point)

OpenStudy (anonymous):

a. 6 in ^3 b. 24 in ^3 c. 36 in ^3 d. 48in ^3 I GOT 24 IN ^3

OpenStudy (anonymous):

@jdoe0001 HELP

OpenStudy (jdoe0001):

\(\bf \textit{volume of a cone}=\cfrac{{\color{red}{ \pi r^2 h}}}{3}\qquad \textit{volume of a cylinder}={\color{red}{ \pi r^2 h}}\) if the cone fits inside a cylinder, that is  thus "h" is the same value for both

OpenStudy (jdoe0001):

\(\bf \textit{volume of a cone}=\cfrac{{\color{red}{ \pi r^2 h}}}{3}\qquad \textit{volume of a cylinder}={\color{red}{ \pi r^2 h}}\\ \quad \\ \textit{volume of a cone}=\cfrac{{\color{red}{ \textit{volume of a cylinder}}}}{3}\\ \quad \\\implies 3\cdot \textit{volume of a cone}={\color{red}{ \textit{volume of a cylinder}}}\)